Fessenden Corporation has accumulated a significant amount of debt as a result of debt-financed acquisitions of other companies. It is currently considering acquiring one of its competitors, Sonar Corporation. Fessenden’s existing debt covenants stipulate that it cannot go beyond a debt to equity ratio of 1.25:1 and a net debt as a percentage of capitalization ratio of 0.90:1. The acquisition of Sonar will cost $88 million. Fessenden’s current level of equity is $460 million and its current level of interest-bearing debt is $621 million. Fessenden has a cash balance of $69 million. It will finance the acquisition with a 10-year bond of $88 million that carries a 5% interest rate sold at par.
Answer #1

ANSWER:

As per the given Question

Part (a)

Debt to equity ratio prior to the proposed transaction = Debt / Equity = 575 / 450 = 1.28 : 1

Net debt as a percentage of capitalization ratio prior to the proposed acquisition = Net debt / total capital = (Debt - cash) / (Debt - Cash + Equity) = (575 - 75) / (575 - 75 + 450) = 500 / 950 = 0.53

Part (b)

Debt to equity ratio after the proposed transaction = Debt / Equity = (575 + 80) / 450 = 1.46 : 1

Net debt as a percentage of capitalization ratio after the proposed acquisition = Net debt / total capital = (Debt - cash) / (Debt - Cash + Equity) = (575 + 80 - 75) / (575 + 80 - 75 + 450) = 580 / 1,030 = 0.53

Fessenden could not acquire Sonar Corporation with the bond issue and still remain in compliance with the existing debt covenants. The first covenant on debt to equity ratio is violated.

thank you for given question.........kindly rate...it helps me a lot